# Learning to utilize image second-order derivative information for crisp edge detection

### v3

@@ Abstract
Index Terms:
Edge detection, Convolutional neural network, Second-order derivative information, Hybrid focal loss function
I Introduction
Edge detection is a fundamental low-level problem in computer vision and image processing, playing a crucial role in a wide range of applications, from autonomous driving to medical imaging. It serves as the cornerstone for many high-level computer vision tasks by identifying the boundaries and structures within images, such as salient object detection [1, 2, 3], objection recognition [4, 5], and image segmentation [6, 7, 8]. With the rapid development of deep learning techniques, numerous researchers aim to employ deep convolutional neural networks (DCNN) for edge detection. Recent studies have proposed several remarkable DCNN-based methods, such as HED [9], RCF [10], and CHRNet [11]. These DCNN-based approaches have achieved state-of-the-art performance on a number of benchmarks including BSDS500 [12] and NYUD-V2 [13], which demonstrate the strong power of DCNN.
@@ I Introduction
However, there is still a major issue that needs to be addressed in edge detection: most modern DCNN-based models [9, 10, 14] are inclined to generate thick and noisy edge maps, such predictions introduce inaccuracies that compromise the spatial precision of high-level tasks, and it is frequently necessary to use a post-processing operation (NMS and morphological thinning) to obtain clean and precise edge maps. This issue can be decomposed into two subsidiary problems: (1) Excessive reliance on DCNN-based autonomous learning, without sufficient prior knowledge of edge information, which is essential for precise edge detection. (2) A highly imbalanced pixel distribution between edges and non-edges in an image, which presents difficulties in pixel classification.
In this work, we attempt to tackle this central issue by addressing the two subsidiary problems. Our goal is to make the network generate one-pixel width edge lines without using the post-processing operation. On the one hand, most DCNN-based methods [9, 10, 14, 15, 16] leverage the powerful feature extraction capabilities of deep learning for edge detection. However, these methods neglect the utilization of prior knowledge regarding image edges, which is typically determined by abrupt changes in pixels. Image second-order derivative captures such changes and it precisely locates the positions of edge pixels by the zero-crossing points. Therefore, leveraging this characteristic can help DCNNs to locate the true edge pixels more accurately. In addition, the mere introduction of second-order derivative information can result in noisy edge maps, as illustrated in Fig. 1 (c). In response to this circumstance, we combine the second-order derivative information with the multi-scale contextual information, developing a second-order derivative-based multi-scale contextual enhancement module (SDMCM). The SDMCM is based on the traditional Laplacian operator [17] which can provide image second-order derivative cues, and dilated convolution with a large receptive field to capture the long-range multi-scale contextual information which can filter out the noise. The proposed SDMCM empowers the network to attain higher precision in true edge pixel localization, consequently generating crisp and clean edge maps.
@@
Furthermore, we employ the conditional parameterized convolution (CondConv) [21] to construct a boundary refinement module (BRM). CondConv can dynamically adjust convolutional kernel parameters according to the contextual information of input feature maps, and the BRM exploits this characteristic to improve the model adaptability to different image scenarios, thereby further suppressing the interference and refining the predicted edge maps. In the end, we propose a LUS-Net which is based on the U-shape architecture for crisp edge detection, and the whole network can be split into three parts: encoder, skip-connections, and decoder. The encoder consists of a lightweight pre-trained model [22, 23, 24] to enable efficient prediction. The SDMCM serves as the skip-connection component. The decoder employs a dense connection to cascade each BRM, allowing the model to learn more diverse and rich feature representations. In addition, we conduct a series of experiments to demonstrate the effectiveness of our method, some instances are shown in Fig. 1 (d).
@@
    We propose a novel hybrid focal loss (HFL) based on the Tversky index and focal loss, which can effectively alleviate the issue of imbalanced pixel distribution, resulting in suppressing misclassified false positive pixels near the true positive pixels.
@@
The paper is structured as follows. Section II presents a review of related work in edge detection. Section III provides an elaboration of the LUS-Net, which involves SDMCM, BRM, and HFL. Section IV presents an analysis of our experiments. We evaluate the crispness of edges, provide detailed descriptions of the ablation study, analyze the function of each component, and compare our approach with recent state-of-the-art algorithms in edge detection. The final section, Section V, summarizes our proposed method and discusses future directions.
II Related work
@@
Traditional methods: Early edge detection methods typically calculate image derivatives to produce edge maps. The Roberts operator [25] is a simple first-order derivative operator that detects edges by computing differences between diagonally adjacent pixels. The Sobel operator [26] is another first-order edge detector that uses two $3\times 3$ kernels to calculate image gradients in horizontal and vertical directions. The Canny [27] detector is a robust algorithm that detects edges by reducing noise, calculating gradients, and applying non-maximum suppression and hysteresis thresholding. The Laplacian detector [17] locates edge pixels by computing the second-order derivative of the image intensity, highlighting regions of rapid intensity change. Over time, researchers have improved edge detection by integrating texture, gradient, and other low-level features. Methods like Pb [28], gPb [12], and SE [29] use a classifier to generate object-level boundaries by utilizing these features. Despite their enhanced performance over derivative-based methods, they still rely on human-designed features and lack semantic information, limiting further improvement.
@@
Deep learning-based methods: Deep learning-based techniques have significantly advanced the area of edge detection and play a crucial role in most state-of-the-art (SOTA) edge detection methods. SOTA edge detectors in recent years mainly adopt convolutional neural networks. These methods achieve significant performance with higher F-scores and some of them even surpass humans on benchmarks such as BSDS500 dataset. HED [9] presents a first end-to-end edge detection architecture, which is built on a fully convolutional VGG-16 network. They generate edge maps by fusing five side-output features with different scales and construct a weighted cross-entropy loss function to address the problem of imbalanced distribution. Based on HED, RCF [10] further utilizes multi-scale features, aggregating all features from different convolutional layers in each stage of VGG-16, improving the ability of the network to capture contextual information, and becoming the first to outperform humans on BSDS500. BDCN [14] proposes an innovative bi-directional cascade network architecture, which leverages information flow from shallow-to-deep and deep-to-shallow to capture multi-scale features within images comprehensively, significantly enhancing the performance. DexiNed [30], inspired by HED and Xception [31], can produce detailed edge maps that are visually appealing, without any pre-training or fine-tuning process. PiDiNet [18] provides a lightweight yet effective solution for edge detection by integrating traditional edge detection operators into vanilla convolution in modern DCNN.
@@
As for precise edge detection, various methods have also been actively explored and have made their contributions. CED [32] focuses on guiding the network to predict sharp edge maps by employing sub-pixel convolution to upsample features. LPCB [33] explains the reason for edge thickness and proposes a new loss function based on the Dice coefficient [34] which enables the network to generate crisp boundaries without requiring post-processing. DSCD [35] aims to produce high-quality edge maps, they introduce a novel loss function inspired by SSIM [36] and build a dense connection [37] hyper-module based on dilated convolution [38] to achieve better performance. DRC [15] presents an innovative network through stacking refinement modules along with an adaptive weighting strategy for the optimal combination between different loss functions, resulting in a satisfactory performance. CATS [39] presents a context-aware tracing strategy which consists of a novel tracing loss and a context-aware fusion block for crisp edge detection. FCL-Net [40] proposes a novel network architecture designed to enhance the accuracy of edge detection through fine-scale corrective learning mechanisms. EDETR [41] is the first Transformer-based [42] edge detector which consists of two stages, they can extract well-defined boundaries of objects and meaningful edges by utilizing contextual information from the entire image, as well as detail local cues.
@@
Our approach is motivated by the above pioneer works [33, 19, 17] to address the problem of edge thickness. We make the network produce crisp and clean edge maps by combining multi-scale contextual and second-order derivative information. Additionally, our method successfully addresses the issue of imbalanced pixel distribution by constructing a hybrid focal loss function, resulting in an excellent performance.
III Methodology
@@
III-A Lightweight pre-trained backbone
Most SOTA edge detection methods adopt a pre-trained VGG [43] or ResNet [44] as a feature extraction backbone and then fine-tune them on the edge detection dataset. However, these backbones, with a large number of parameters, demand an expensive computational cost. To achieve efficient edge detection, we adopt and test three lightweight pre-trained backbones: MobileNetV2 [22], ShuffleNetV2 [23], and EfficientNetV2 [24]. These lightweight backbones have demonstrated comparable performance to VGG or ResNet networks while providing efficient inference capabilities. Specifically, these models employ a stride of 2 in the initial convolutional layer, resulting in excessive downsampling. This leads to feature maps with significantly reduced spatial resolution, damaging the quality of the generated contours. Therefore, we modify the stride from 2 to 1. Additionally, we remove the classifier head to make the backbone network suitable for edge detection, which simultaneously reduces the network parameters.
III-B Second-order derivative-based multi-scale contextual enhancement module
@@
In the multi-scale contextual path, we first employ a channel compression operation $\mathcal{C}(\cdot)$ which is a $1\times 1$ convolution to compress the input feature channel. This operation effectively prevents overfitting while simultaneously reducing the number of parameters. The compression ratio $r\in\left\{\frac{1}{2},\frac{1}{4},\frac{1}{8}\right\}$, we find that changes in compression rate have an impact on the crispness of the predicted edge maps (Section IV provides details). After the channel compression, we construct four parallel branches $\left\{\mathcal{B}_{1}(\cdot),\mathcal{B}_{2}(\cdot),\mathcal{B}_{3}(\cdot),\mathcal{B}_{4}(\cdot)\right\}$ to produce multi-scale contextual features, and each branch consists of several $3\times 3$ dilated Conv-ReLU sequences. Introducing dilated convolution using a fixed dilated rate can broaden the receptive field, however, it also produces fixed feature intervals, which leads to blurred and discontinuous edges. Therefore, we stack a series of dilated convolutions with different dilated rates $d\in\left\{1,2,3\right\}$, allowing for filling intervals in the features. In this way, we can utilize all the features in continuous convolution kernels, without producing unused features or fixed intervals, and capturing long-range semantic information by expanding the receptive field. In the end, we combine the features from four branches by element-wise sum to obtain the mixed features which contain rich multi-scale contextual information.
@@
$Y=$ $\mathcal{S}\left[\mathcal{B}_{1}\left(\mathcal{C}(X)\right)+\mathcal{B}_{2}\left(\mathcal{C}(X)\right)+\mathcal{B}_{3}\left(\mathcal{C}(X)\right)+\mathcal{B}_{4}\left(\mathcal{C}(X)\right)\right.$ (3)
$\left.+\mathcal{C}(X)\right]$
@@
III-C Boundary refinement module
@@
III-D Hybrid focal loss
As a pixel-level binary classification task, edge detection encounters a significant challenge of imbalanced class distribution. The edge pixels typically comprise only about 10% of an image. Such an imbalanced issue causes great difficulties in training the network. Current mainstream methods use the weighted cross-entropy loss to tackle this, but it often results in thick boundaries [33]. To address the issue of imbalanced class distribution, we propose a hybrid focal loss function.
Imbalanced class distribution is a common issue in object detection and image segmentation. To address this, we draw on solutions from these two fields. Specifically, the focal loss [20] in object detection is effective in scenarios with extreme class imbalance and the need for fine detail detection, the Tversky index [19] in image segmentation shines in situations requiring specific trade-off between false positives and false negatives. These characteristics are crucial for edge detection.
The focal loss aims to tackle the issue of imbalanced data sample distribution, and we leverage its idea to address the problem of imbalanced pixel distribution in an image. In pixel-level binary classification, it is well-known that the standard cross-entropy loss function is used to evaluate the proximity of the predictions to the labels. However, in imbalanced pixel distribution, the loss is dominated by easily classified pixels from the majority class, neglecting the harder-to-learn minority class. In this situation, we introduce a focusing parameter $\gamma\geq 0$ into cross-entropy to downweigh the loss for easy (well-classified) pixels and focus training on hard (misclassified) pixels. This ensures the model pays more attention to the minority class and improves its ability to correctly classify these challenging cases, alleviating the imbalanced pixel distribution. The focal loss can be written as:
@@ III-D Hybrid focal loss
where $p_{i}$ and $g_{i}$ represent the value of $i$-th pixel on a predicted edge map and its corresponding groundtruth image, respectively. $N$ is the total number of pixels in an image. $(1-p_{i})^{\gamma}$ is a modulating factor and $\alpha$ is a balance factor for positive and negative pixels.
The focal loss function is based on pixel error. When dealing with class imbalance, using such a strategy only will result in correct classification but insufficient accuracy, which leads to generating thick edge maps. To tackle the issue of thickness, we introduce the image-level information and construct a focal Tversky loss function based on the Tversky index. The Tversky index can be written as follows:
@@
where P indicates the predicted edge maps and G indicates its corresponding labels. $\alpha$ and $\beta$ control the weight of false negatives (FNs) and false positives (FPs), respectively. By adjusting the value of $\alpha$ and $\beta$, we can make a trade-off between FN pixels and FP pixels.
Inspired by the focal loss modification of the cross-entropy loss, we adopt a similar strategy and introduce $\gamma$ into the Tversky index. The focal Tversky loss function can be defined as: {strip}
@@
where $p_{i}(1-g_{i})$ and $(1-p_{i})g_{i}$ represent FPs and FNs, respectively. $\alpha+\beta=1$. $\gamma$ is the focusing parameter to control the easy-training non-edge pixels and hard-training edge pixels, and $C$ is a constant number to prevent the numerator/denominator from becoming 0. In this equation, we set $\alpha=0.3$, $\beta=0.7$, because the FN pixels are more important in edge detection. In addition, $\gamma=0.75$ which places greater emphasis on low-accuracy predictions that have been misclassified, and $C=1\times 10^{-7}$. In the training process, the focal Tversky loss function maximum the value of $\sum_{i=1}^{N}p_{i}g_{i}$ which represents the TP pixels and thereby enables the model to locate the true edge pixels more accurately.
@@
where we set $\lambda=0.001$ to balance the weight between focal loss and focal Tversky loss. The hybrid focal loss function constrains the network from both pixel-level and image-level information. By leveraging the focal loss, the model concentrates on pixels that are challenging to classify while reducing the influence of well-classified pixels, resulting in alleviating the issue of imbalanced distribution. Simultaneously, by integrating the focal Tversky loss, the model can locate the true positive edge pixels more precisely and generate crisp boundaries by utilizing the global image-level information. Therefore, our hybrid focal loss function can tackle both class imbalance issues and thickness problems, consequently improving the performance of the model.
IV Experiments
In this section, we provide a comprehensive account of the implementation details, encompassing hyperparameters, the adopted datasets, and their augmentation strategy. We then introduce the two evaluation methods employed in this work, followed by a series of ablation experiments on our approach. Finally, we compare our proposed method with some state-of-the-art (SOTA) edge detection algorithms and demonstrate its superiority.
IV-A Implementation details

### v4

@@ Abstract
keywords:
[…]
1 Introduction
Edge detection is a fundamental problem in computer vision and image processing, playing a crucial role in a wide range of applications, from autonomous driving to medical imaging. It serves as the cornerstone for many high-level computer vision tasks by identifying the boundaries and structures within images, such as salient object detection [1, 2, 3, 4, 5, 6], objection recognition [7, 8, 9, 10, 11], image segmentation [12, 13, 14, 15, 16, 17, 18, 19, 20], and object tracking [21, 22, 23]. With the rapid development of deep learning techniques, numerous researchers aim to employ Deep Convolutional Neural Networks (DCNNs) for edge detection. Recent studies have proposed several remarkable DCNN-based methods, such as HED [24], RCF [25], and CHRNet [26]. These DCNN-based approaches have achieved state-of-the-art performance on a number of benchmarks including BSDS500 [27] and NYUD-V2 [28], which demonstrate the strong power of DCNN.
@@ 1 Introduction
However, there is still a major issue that needs to be addressed in edge detection: most modern DCNN-based models [24, 25, 29] are inclined to generate thick and noisy edge maps, such predictions introduce inaccuracies that compromise the spatial precision of high-level tasks, and it is frequently necessary to use a post-processing operation (NMS and morphological thinning) to obtain clean and precise edge maps. This issue can be decomposed into two subsidiary problems: (1) Excessive reliance on DCNN-based autonomous learning, without sufficient prior knowledge of edge information, which is essential for precise edge detection; (2) A highly imbalanced pixel distribution between edges and non-edges in an image, which presents difficulties in pixel classification.
In this work, we attempt to tackle this central issue by addressing the two subsidiary problems. Our goal is to make the network generate one-pixel width edge lines without using the post-processing operation. On the one hand, most DCNN-based methods [24, 25, 29, 30, 31] leverage the powerful feature extraction capabilities of deep learning for edge detection. However, these methods neglect the utilization of prior knowledge regarding image edges, which are typically determined by abrupt changes in pixels. Image second-order derivative captures such changes and it precisely locates the positions of edge pixels by the zero-crossing points. Therefore, leveraging this characteristic can help DCNNs to locate the true edge pixels more accurately. In addition, the mere introduction of second-order derivative information can result in noisy edge maps, as illustrated in Fig. 1 (c). In response to this circumstance, we combine the second-order derivative information with the multi-scale contextual information, developing a Second-order Derivative-based Multi-scale Contextual Enhancement Module (SDMCM). The SDMCM is based on the traditional Laplacian operator [32] which can provide image second-order derivative cues, and dilated convolution with a large receptive field to capture the long-range multi-scale contextual information which can filter out the noise. The proposed SDMCM empowers the network to attain higher precision in true edge pixel localization, consequently generating crisp and clean edge maps.
@@
Furthermore, we employ the Conditional parameterized Convolution (CondConv) [36] to construct a Boundary Refinement Module (BRM). CondConv can dynamically adjust convolutional kernel parameters according to the contextual information of input feature maps, and BRM exploits this characteristic to improve the model adaptability to different image scenarios, thereby further suppressing the interference and refining the predicted edge maps. In the end, we propose a LUS-Net which is based on the U-shape architecture for crisp edge detection, and the whole network can be split into three parts: encoder, skip-connections, and decoder. The encoder consists of a lightweight pre-trained model [37, 38, 39] to enable efficient prediction. The SDMCM serves as the skip-connection component. The decoder employs a dense connection to cascade each BRM, allowing the model to learn more diverse and rich feature representations. In addition, we conduct a series of experiments to demonstrate the effectiveness of our method, some instances are shown in Fig. 1 (d).
@@
    We propose a novel Hybrid Focal Loss (HFL) based on Tversky index and focal loss, which can effectively alleviate the issue of imbalanced pixel distribution, resulting in suppressing misclassified false positive pixels near the true positive pixels.
@@
The paper is structured as follows. Section 2 presents a review of related work in edge detection. Section 3 provides an elaboration of the LUS-Net, which involves SDMCM, BRM, and HFL. Section 4 presents an analysis of our experiments. We evaluate the crispness of edges, provide detailed descriptions of the ablation study, analyze the function of each component, and compare our approach with recent state-of-the-art algorithms in edge detection. The final section, Section 5, summarizes our proposed method and discusses future directions.
2 Related work
@@
Traditional methods: Early edge detection methods typically calculate image derivatives to produce edge maps. The Roberts operator [40] was a simple first-order derivative operator that detected edges by computing differences between diagonally adjacent pixels. The Sobel operator [41] was another first-order edge detector that used two $3\times 3$ kernels to calculate image gradients in horizontal and vertical directions. The Canny detector [42] was a robust algorithm that detected edges by reducing noise, calculating gradients, and applying non-maximum suppression and hysteresis thresholding. The Laplacian detector [32] located edge pixels by computing the second-order derivative of the image intensity, highlighting regions of rapid intensity change. Over time, some researchers improved edge detection by integrating texture, gradient, and other low-level features. Methods like Pb [43], gPb [27], and SE [44] used a classifier to generate object-level boundaries by utilizing these features. Despite their enhanced performance over derivative-based methods, they still relied on human-designed features and lacked semantic information, limiting further improvement.
@@
Deep learning-based methods: Deep learning-based techniques significantly advanced the area of edge detection and played a crucial role in most state-of-the-art (SOTA) edge detection methods. SOTA edge detectors in recent years have mainly adopted convolutional neural networks. These methods achieved significant performance with higher F-scores and some of them even surpassed humans on benchmarks such as BSDS500 dataset. Xie et al. [24] presented the first end-to-end edge detection architecture named Holistically-nested Edge Detection (HED), which was built on a fully convolutional VGG-16 network. They generated edge maps by fusing five side-output features with different scales and constructed a weighted cross-entropy loss function to address the problem of imbalanced pixel distribution. Based on HED, Liu et al. [25] further utilized multi-scale features by aggregating all features from different convolutional layers in each stage of VGG-16. This improved the network’s ability to capture contextual information, making it the first to outperform humans on BSDS500. He et al. [29] proposed an innovative bi-directional cascade network architecture, which leveraged information flow from shallow-to-deep and deep-to-shallow to capture multi-scale features within images comprehensively, significantly enhancing the performance. Soria et al. [45] proposed DexiNed, a network inspired by HED and Xception [46], which could produce detailed edge maps that were visually appealing without any pre-training or fine-tuning process. Su et al. [33] provided a lightweight yet effective solution for edge detection by integrating traditional edge detection operators into vanilla convolution in modern DCNN.
@@
Regarding precise edge detection, various methods have contributed significantly to the field. Wang et al. [47] focused on guiding the network to predict sharp edge maps by employing sub-pixel convolution to upsample features. Deng et al. [48] explained the reason for edge thickness and proposed a new loss function based on the Dice coefficient [49] which enables the network to generate crisp boundaries without requiring post-processing. Deng et al. [50] proposed DSCD to produce high-quality edge maps by introducing a novel loss function inspired by SSIM [51] and built a dense connection [52] hyper-module based on dilated convolution [53]. Cao et al. [30] presented an innovative network by stacking refinement modules and implementing an adaptive weighting strategy for the optimal combination between different loss functions, resulting in a competitive performance. Huan et al. [54] presented a context-aware tracing strategy using a novel tracing loss and a context-aware fusion block for crisp edge detection. Xuan et al. [55] proposed a novel network architecture designed to enhance the accuracy of edge detection through fine-scale corrective learning mechanisms. Liu et al. [56] proposed GCB-Net, a U-shape encoder-decoder architecture with feature enhancement and fusion modules, which could generate crisp and clean edge maps through multi-scale features and a novel mixed loss function based on Tversky index. Pu et al. [57] EDTER, the first Transformer-based [58] edge detector with a two-stage architecture. The model could extract well-defined boundaries of objects and meaningful edges by utilizing contextual information from the entire image, as well as detail local cues. Zhou et al. [59] designed an edge granularity network to estimate edge granularity from annotations and integrated it into both spatial and frequency domains of feature maps, enabling the generation of diverse edge predictions ranging from coarse object contours to fine textures. Cetinkaya et al. [60] constructed a ranking-based approach to address both class imbalance and label uncertainty in edge detection through ranking positive pixels over negatives and promoting high confidence edge pixels.
@@
Despite their remarkable performance, existing methods still fall short in adequately addressing the challenge of edge thickness. Our approach is motivated by these pioneering works [48, 34, 32, 56] to address the problem of edge thickness. Through the combination of multi-scale contextual and second-order derivative information, we enable the network to produce crisp and clean edge maps. Additionally, our method successfully addresses the issue of imbalanced pixel distribution by constructing a hybrid focal loss function, resulting in excellent performance.
3 Methodology
@@
3.1 Lightweight pre-trained backbone
Most SOTA edge detection methods adopt a pre-trained VGG [61] or ResNet [62] as a feature extraction backbone and then fine-tune them on the edge detection dataset. However, these backbones, with a large number of parameters, demand an expensive computational cost. To achieve efficient edge detection, we adopt and test three lightweight pre-trained backbones: MobileNetV2 [37], ShuffleNetV2 [38], and EfficientNetV2 [39]. These lightweight backbones have demonstrated comparable performance to VGG or ResNet networks while providing efficient inference capabilities. Specifically, these models employ a stride of two in the initial convolutional layer, resulting in excessive downsampling. This leads to feature maps with significantly reduced spatial resolution, damaging the quality of the generated contours. Therefore, we modify the stride from two to one. Additionally, we remove the classifier head to make the backbone network suitable for edge detection, which simultaneously reduces the network parameters.
3.2 Second-order derivative-based multi-scale contextual enhancement module
@@
In the multi-scale contextual path, we first employ a channel compression operation $\mathcal{C}(\cdot)$ which is a $1\times 1$ convolution to compress the input feature channel. This operation effectively prevents overfitting while simultaneously reducing the number of parameters. The compression ratio $r\in\left\{\frac{1}{2},\frac{1}{4},\frac{1}{8}\right\}$, we find that changes in compression rate have an impact on the crispness of the predicted edge maps (Section 4 provides details). After the channel compression, we construct four parallel branches $\left\{\mathcal{B}_{1}(\cdot),\mathcal{B}_{2}(\cdot),\mathcal{B}_{3}(\cdot),\mathcal{B}_{4}(\cdot)\right\}$ to produce multi-scale contextual features, and each branch consists of several $3\times 3$ dilated Conv-ReLU sequences. Introducing dilated convolution using a fixed dilated rate can broaden the receptive field, however, it also produces fixed feature intervals, which leads to blurred and discontinuous edges. Therefore, we stack a series of dilated convolutions with different dilated rates $d\in\left\{1,2,3\right\}$, allowing for filling intervals in the features. In this way, we can utilize all the features in continuous convolution kernels, without producing unused features or fixed intervals, and capturing long-range semantic information by expanding the receptive field. In the end, we combine the features from four branches by element-wise sum to obtain the mixed features which contain rich multi-scale contextual information.
@@
$Y=$ $\mathcal{S}\left[\mathcal{B}_{1}\left(\mathcal{C}(X)\right)+\mathcal{B}_{2}\left(\mathcal{C}(X)\right)+\mathcal{B}_{3}\left(\mathcal{C}(X)\right)+\mathcal{B}_{4}\left(\mathcal{C}(X)\right)+\mathcal{C}(X)\right]$ (3)
@@
3.3 Boundary refinement module
@@
3.4 Hybrid focal loss
As a basic computer vision task, edge detection encounters a significant challenge of imbalanced class distribution, with edge pixels typically comprising only about 10% of an image. While current mainstream methods use weighted cross-entropy loss to tackle this issue, they often result in thick boundaries as such strategy lacks global structural information [48]. Therefore, edge detection demands more complete solutions beyond simple class balancing, as it requires both pixel-level precision and global structural information. Inspired by the focal loss [35] and Tversky index [34], we propose a hybrid approach that specifically addresses these unique challenges. Our key insight is that effective edge detection requires simultaneous optimization of both local pixel-level accuracy and global structural information, which neither existing loss function alone can achieve.
The focal loss aims to tackle the issue of imbalanced data sample distribution, and we adopt this concept to tackle the imbalanced pixel distribution in an image. While standard cross-entropy has been widely used in pixel-level binary classification to measure prediction accuracy, it exhibits significant limitations when dealing with imbalanced distributions. Specifically, the loss function becomes predominantly influenced by easily classified majority-class pixels, leading to insufficient learning of the more challenging minority-class pixels. Therefore, a focusing parameter $\gamma\geq 0$ is introduced into cross-entropy to downweigh the loss for well-classified pixels and focus training on misclassified pixels. The focal loss can be written as:
@@ 3.4 Hybrid focal loss
where $p_{i}$ and $g_{i}$ represent the value of $i$-th pixel on a predicted edge map and its corresponding groundtruth image, respectively. $N$ is the total number of pixels in an image. $(1-p_{i})^{\gamma}$ is a modulating factor and $\alpha$ is a balance factor for positive and negative pixels. When $\gamma\geq 0$, the modulating factor $(1-p_{i})^{\gamma}$ produces an adaptive scaling effect. When $p_{i}$ approaches 1, the factor reduces the loss for well-classified pixels. Conversely, when $p_{i}$ is small, the factor maintains high loss for misclassified pixels. This adaptive modulation mechanism ensures that the model pays more attention to the minority class and improves its ability to correctly classify these challenging cases, providing a principled solution to the imbalanced pixel distribution.
Although focal loss can address class imbalance and achieve correct pixel classification, it primarily operates on local pixel relationships, neglecting global structural information. This limitation impairs its ability to suppress false edge pixels around true edge pixels, often resulting in thick edge maps. To address this issue, we introduce global structural information through the Tversky index, which is formally defined as follows:
@@
where P indicates the predicted edge maps and G indicates its corresponding labels. $\alpha$ and $\beta$ control the weight of false negatives (FNs) and false positives (FPs), respectively. The sum of these two parameters is set to 1.
Inspired by the focal loss modification of the cross-entropy loss, we adopt a similar strategy and introduce $\gamma$ into the reciprocal of Tversky index. The focal Tversky loss function can be defined as:
@@
where $p_{i}(1-g_{i})$ and $(1-p_{i})g_{i}$ represent FPs and FNs, respectively. $C=1\times 10^{-7}$ is a constant number to prevent the numerator/denominator from being 0.
Here, $\gamma=0.75$ serves as a focusing parameter. Specifically, when the $\frac{1}{Tversky}\gg$ 1, suggesting predictions with errors, $\gamma$ reduces the loss significantly. Conversely, when the $\frac{1}{Tversky}$ is close to 1, indicating accurate pixel classification, the loss shows high fluctuation. Therefore, during early training, as the network lacks global structural information, it emphasizes pixel-level predictions. The reduction effect of $\gamma$ enables the focal loss to dominate the primary learning process. As training process to later stages, most pixels are well-classified, and the $\frac{1}{Tversky}$ approaches 1. As evidenced by Fig. 6, the loss function exhibits steeper gradients near $\frac{1}{Tversky}$ close to 1, focal Tversky loss dominates the learning process at this point, enabling the model to focus more attention on these well-classified pixels for further refinement by utilizing the global structural information.
Figure 6: Visualization of focal Tversky loss curves under different $\gamma$.
@@
where we set $\lambda=0.001$ to balance the weight between focal loss and focal Tversky loss. The focal loss operates at the pixel level, using adaptive scaling effect to address class imbalance, while the focal Tversky loss operates at the image level, leveraging global structural information for accurate pixel location. Therefore, our hybrid focal loss function can tackle both class imbalance issues and thickness problems, consequently improving the performance of the model.
4 Experiments
In this section, we provide a comprehensive account of the implementation details, encompassing hyperparameters, the adopted datasets, and their augmentation strategy. We then introduce the two evaluation methods employed in this work, followed by a series of ablation experiments on our approach. Finally, we compare our proposed method with some SOTA edge detection algorithms and demonstrate its superiority.
4.1 Implementation details